# AnyTop: Character Animation Diffusion with Any Topology

### v1

@@ 6.2. Implementation details
We use $T=100$ diffusion steps, $L=4$ STT layers, and latent dimension $F=128$. We train the model using a single NVIDIA RTX A6000 GPU for 24 hours. Inference runs on an NVIDIA GeForce RTX 2080 Ti GPU. More implementation details can be found in Appendix A .
@@ 6.4. Baselines
To the best of our knowledge, no current works address such a diverse range of skeletal structures within a single model. Hence, we compare AnyTop to adaptations of two baselines. The first is MDM (Tevet et al., 2023), originally designed for a single humanoid skeleton. MDM uses per-frame embedding, so to match its representation format, we concatenate all joint features for each character, and pad them to a length of $J\times D$. For fairness, we also concatenate the vectorized rest-pose embedding $\mathcal{P}_{\mathcal{S}}$ along the temporal axis as frame 0. Since MDM accepts textual conditions, we use the skeleton’s name (e.g., Cat, Dragon) as the input text. Additionally, since MDM’s original configuration was designed for a dataset 14 times larger than ours (Petrovich et al., 2022b), we reduced its latent dimension size to mitigate overfit.
@@ 6.5. Quantitative Results
Table 2 shows a quantitative comparison of AnyTop and the baselines. AnyTop outperforms MDM in all categories and SinMDM in all but coverage, which is expected since SinMDM is trained separately for each skeleton. Note the significant gap in diversity metrics, where the table shows AnyTop generalizes well, while the others struggle to do so. We also report the models’ parameter count, showing ours uses fewer parameters, enabling lower computation and faster inference. In Appendix C , we provide a comparison with the baselines on the data subsets, demonstrating our model’s superiority on these as well.
@@ 6.6. Qualitative Results
Unseen skeleton
We present two unseen skeleton motions. One is a komodo dragon, generated by the Bipeds model. The second is a Cat, generated by a model trained on Quadrupeds, excluding the cat. Figure 8 and our supp. video demonstrate AnyTop generalizes well to unseen skeletons, while adapted MDM under the same settings generates static and jittery motions.
@@ 6.7. Ablation
In table 3, we explore three key components of AnyTop’s architecture. First, the results confirm that without access to topological information, the model struggles to prioritize joints based on their hierarchical relations. Omitting the incorporation of $\mathcal{D}$ and $\mathcal{R}$ leads to degradation in all metrics. Next, excluding the rest pose $\mathcal{P}_{\mathcal{S}}$ produces inferior results, reinforcing the idea that $\mathcal{P}_{\mathcal{S}}$ encodes vital information about joint offsets and bone lengths. Lastly, we examine cross-skeletal prior sharing via the addition of joint name embeddings. While cross-gen improves motion diversity, it introduces a tradeoff, as generated motions may exhibit motifs absent in the skeleton’s ground truth, reducing coverage. Results show that removing joint name embeddings increases coverage but severely sacrifices diversity and cross-skeleton generalization.
@@ 7. Conclusion, Limitations and Future Work
One limitation of our method stems from imperfections in the input data. Despite our cleaning procedure, certain data artifacts remain unresolved. Another limitation is that our data augmentation process is computationally expensive with $O(J^{2})$ complexity.
@@ 8. Acknowledgements
In Appendix A, we provide implementation details of AnyTop, and in Appendix B, we elaborate on our preprocessing and augmentation pipelines. Finally, in Appendix C and present additional quantitative results beyond those in the main paper.
Appendix A Implementation Details
The maximum topological distance we allow in $\mathcal{D}$ is $d_{max}=6$, and our Temporal Attention is applied on temporal windows of length $W=31$. For our model inputs, we allow maximum number of joints $J=143$. During training, we use cropped sequences of $N=40$ frames. To enable our model handle higher frame positions and generate longer sequences, we incorporate positional encoding relative to the cropping index. For training, we used batch size of 16 when training on the entire dataset, and a batch size of 8 to train in the data subsets.
Appendix B Data
@@
In addition to the alignment procedure, we also extract relevant information from the skeletons and motion data. First, we use foot joints labels to generate foot-contact indicators for each frame, which are concatenated with the motion features. Next, we compute the mean and standard deviation for each skeleton’s frames and use these statistics to normalize the motions before feeding them into the model during training.
@@
Skeletal Augmentation exposes our model to a wider variety of skeletons, as described in the Method section of the main paper. Next, we further elaborate about this process. The first augmentation we apply is joint removal, which randomly removes up to 30% of the joints from the skeleton, where feet joints are never removed to maintain physical correctness. For efficiency consideration, we exclude joints with more than a single child from the removal procedure. The second augmentation is joint addition, which introduces a new joint at the midpoint of a randomly selected edge. After removing or adding joints to the skeleton, we update $\mathcal{R}_{\mathcal{S}}$, $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{N}_{\mathcal{S}}$ accordingly. Note that updating $\mathcal{D}_{\mathcal{S}}$ is computationally expensive with a complexity of $0(J^{2})$, as it requires recomputing the path between each pair of joints in the DAG.
Appendix C Comparison with Baselines on Subset Models
@@
We provide a quantitative evaluation of AnyTop trained on the data subsets defined in the Experiments section of the main paper. To maintain fairness in comparison, we train the adapted MDM baseline separately for each subset. Since SinMDM is independently trained for each skeleton, no additional adjustments are needed. Each model is evaluated using the corresponding skeletons from our benchmark that match the relevant data subset. The results, shown in table 4, indicate that AnyTop achieves the optimal coverage-diversity tradeoff compared to all other baselines presented.

### v2

@@ 6.2. Implementation details
We use $T=100$ diffusion steps, $L=4$ STT layers, and latent dimension $F=128$. We train the model using a single NVIDIA RTX A6000 GPU for 24 hours. Inference runs on an NVIDIA GeForce RTX 2080 Ti GPU. More implementation details can be found in Appendix B.
@@ 6.4. Baselines
To the best of our knowledge, no current works address such a diverse range of skeletal structures within a single model. Hence, we compare AnyTop to adaptations of two baselines. The first is MDM (Tevet et al., 2023), originally designed for a single humanoid skeleton. MDM uses per-frame embedding, so to match its representation format, we concatenate all joint features for each character, and pad them to a length of $J\times D$. For fairness, we also concatenate the vectorized rest-pose embedding $\mathcal{P}_{\mathcal{S}}$ along the temporal axis as frame 0. Since MDM accepts textual conditions, we use the skeleton’s name (e.g., Cat, Dragon) as the input text. Additionally, since MDM’s original configuration is designed for a dataset 14 times larger than ours (Guo et al., 2022), we reduce its latent dimension size to mitigate overfit.
@@ 6.5. Quantitative Results
Table 3 shows a quantitative comparison of AnyTop and the baselines. AnyTop outperforms MDM in all categories and SinMDM in all but coverage, which is expected since SinMDM is trained separately for each skeleton. Note the significant gap in diversity metrics, where the table shows AnyTop generalizes well, while the others struggle to do so. We also report the models’ parameter count, showing ours uses fewer parameters, enabling lower computation and faster inference. In Appendix D, we provide a comparison with the baselines on the data subsets, demonstrating our model’s superiority on these as well.
@@ 6.6. Qualitative Results
Unseen skeleton
We present two unseen skeleton motions. One is a Komodo Dragon, generated by the Bipeds model. The second is a Cat, generated by a model trained on Quadrupeds, excluding the cat. Figure 8 and our supp. video demonstrate that AnyTop generalizes well to unseen skeletons, while adapted MDM under the same settings generates static and jittery motions.
@@ 6.7. Ablation
In Tab. 4, we explore three key components of AnyTop’s architecture. First, the results confirm that without access to topological information, the model struggles to prioritize joints based on their hierarchical relations. Omitting the incorporation of $\mathcal{D}$ and $\mathcal{R}$ leads to degradation in all metrics. Next, excluding the rest pose $\mathcal{P}_{\mathcal{S}}$ produces inferior results, reinforcing the idea that $\mathcal{P}_{\mathcal{S}}$ encodes vital information about joint offsets and bone lengths. Lastly, we examine cross-skeletal textual prior sharing via the addition of joint name embeddings. Incorporating joint names can be interpreted as balancing between in-skeleton and cross-skeleton generalization. Our ablation shows that excluding joint names (encouraging in-gen) increases coverage, whereas incorporating them (encouraging cross-gen) increases diversity. By employing joint names, we enhance spatial correspondence, cross-generalization and unseen-generalization, acknowledging a minor decrease in coverage as an acceptable cost.
@@ 7. Conclusion, Limitations and Future Work
One limitation of our method stems from imperfections in the input data. Despite our cleaning procedure, certain data artifacts remain unresolved, as detailed in Appendix C. Another limitation is that our data augmentation process is computationally expensive with $O(J^{2})$ complexity.
@@ Acknowledgements.
In Appendix B, we provide implementation details of AnyTop, and in Appendix C, we elaborate on our preprocessing and augmentation pipelines. Finally, in Appendix D, we present additional quantitative results beyond those in the main paper.
Appendix A Related Work - Additional Details
Table 5. Methods categorized by supported skeletal variability. Each column features representative research works that support the type in its header, with full definitions of these categories provided in the main paper.
Single Isomorphic Homeomorphic Non-homeomorphic
Text2Motion (Guo et al., 2022) R2ET (Zhang et al., 2023b) SAN (Aberman et al., 2020) MoMa (Martinelli et al., 2024)
ACTOR (Petrovich et al., 2021) (Villegas et al., 2021) SAME (Lee et al., 2023) WalkTheDog (Li et al., 2024)
GMD (Karunratanakul et al., 2023) SMPL (Loper et al., 2015) SMRNet (Zhang et al., 2024e)
MAS (Kapon et al., 2023) HUMOS (Tripathi et al., 2025) CAR (Cao and Yang, 2024)
MoFusion (Dabral et al., 2023) SCT (Jang et al., 2024) BMM (Studer et al., 2024)
MDM (Tevet et al., 2023) TEMOS (Petrovich et al., 2022) DragPoser (Ponton et al., 2024)
(Holden et al., 2016) SMAL (Zuffi et al., 2017) (Zhang et al., 2024b)
MotionDiffuse (Zhang et al., 2024a) OmniMotionGPT (Yang et al., 2023)
PriorMDM (Shafir et al., 2024) BARC (Rueegg et al., 2023)
MoDi (Raab et al., 2023)
MoMask (Guo et al., 2024)
Tab. 5 categorizes the methods listed in the main paper according to their supported skeletal variability.
Appendix B Implementation Details
Table 6. Ablation on the value of $d_{max}$.
$d_{max}$ Coverage $\uparrow$ Local Div. $\uparrow$ Inter Div. $\uparrow$ Intra Div. Diff. $\downarrow$
1 71.8 0.25 0.27 0.11
5 (Ours) 80.5 0.25 0.31 0.11
10 79.3 0.23 0.29 0.12
The maximum topological distance we allow in $\mathcal{D}$ is $d_{max}=5$, which is the closest to the dataset’s average kinematic chain length (4.65). In Tab. 6, we show that indeed $d_{max}=5$ induces better scores. Our Temporal Attention is applied on temporal windows of length $W=31$, and for our model inputs, we allow a maximum number of joints $J=143$. During training, we use cropped sequences of $N=40$ frames. To enable our model to handle higher frame positions and generate longer sequences, we incorporate positional encoding relative to the cropping index. For training, we use a batch size of 16.
Appendix C Data
@@
In addition to the alignment procedure, we also extract relevant information from the skeletons and motion data. First, we use foot joints labels to generate foot-contact indicators for each frame, which are concatenated with the motion features.
Next, we apply a normalization procedure. Note that Truebones aggregates data from multiple sources, which leads to skeletons of inconsistent sizes, irrespective of their real-world proportions. This procedure involves computing the mean and standard deviation for each skeleton across all frames in all its motions and using these statistics to normalize the motion data before training the model.
Despite our thorough cleaning procedure, several data artifacts remain unresolved in the ground-truth (GT) motions: (1) inherent foot sliding effects (as in the walking chicken GT example in our supplementary video); (2) artificial bones connecting characters to the ground; (3) anatomical inconsistencies, such as right foreleg emerging from left shoulder and left foreleg from right shoulder; (4) extraneous bones protruding from various body joints.
@@
Skeletal Augmentation exposes our model to a wider variety of skeletons, as described in the Method section of the main paper. Next, we further elaborate about this process. The first augmentation we apply is joint removal, which randomly removes up to 30% of the joints from the skeleton, where feet joints are never removed to maintain physical correctness. For efficiency, we exclude joints with multiple children from the removal procedure. We remove only those chains that terminate with end effectors, with the exception of feet. The second augmentation is joint addition, which introduces a new joint at the midpoint of a randomly selected edge. After removing or adding joints to the skeleton, we update $\mathcal{R}_{\mathcal{S}}$, $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{N}_{\mathcal{S}}$ accordingly. Note that updating $\mathcal{D}_{\mathcal{S}}$ is computationally expensive with a complexity of $0(J^{2})$, as it requires recomputing the path between each pair of joints in the DAG.
Appendix D Comparison with Baselines on Subset Models
@@
We provide a quantitative evaluation of AnyTop trained on the data subsets defined in the Experiments section of the main paper. To maintain fairness in comparison, we train the adapted MDM baseline separately for each subset. Since SinMDM is independently trained for each skeleton, no additional adjustments are needed. Each model is evaluated using the corresponding skeletons from our benchmark that match the relevant data subset. The results, shown in Tab. 7, indicate that AnyTop achieves the optimal coverage-diversity tradeoff compared to all other baselines presented.